# HAMLET: Hyperadaptive Agent-based Modeling for Live Embodied Theatrics

### v1

@@ Introduction
To further demonstrate the working mechanism of the HAMLET framework and each of its component designs, we select some real cases as shown in Table 6 in Appendix C. In Case 1, the narrator first considers the situation before responding to make a reasonable judgment about the AI actor’s input. Our framework also supports users taking on a certain character role (Cases 2, 3, and 4), where the narrator, transfer, and advancer can handle various circumstances (such as non-existent props, irresponsible actions, and stubborn choices) to ensure the drama runs smoothly. On the other hand, the planner is responsible for multi-trajectory design and review: the same goal can be met by different beat trajectories as long as the process is reasonable (Cases 5 and 6).
Annotation Statement
This work relies on human annotation, which is critical for both the training and evaluation of our proposed models. The annotation of PAD and HAMLETJudge datasets was conducted by a team of five human experts, each with at least one year of professional experience in relevant fields such as creative writing or drama. To ensure high-quality annotations, we provided clear and task-specific instructions to all labelers. In addition, the datasets underwent rigorous safety checks to ensure no personally identifiable information or potentially problematic content is available. The datasets will be released under the CC BY-NC 4.0 license. The labeling instructions can be found in Appendix E and F.
Conclusion
@@
In this paper, we introduced HAMLET, a multi-agent framework designed to address key challenges in AI-driven theatre. HAMLET generates a guiding narrative blueprint from a simple topic and provides agents with a perceive and decide module, enabling autonomous thinking and physical environmental interaction. To assess performance, we established a comprehensive evaluation method and a leaderboard, where our specialized critic model, HAMLETJudge, achieved top-ranking results. Extensive experiments show that our approach creates expressive and coherent theatrical experiences, paving a new path toward autonomous and immersive interactive drama.
Appendix A A. Detailed information of 100 Cases Selected for Public Dataset
@@
This appendix details the public dataset of 100 cases in Table 4. The dataset comprises three parts: 1) excerpts from 25 Chinese literary works, 2) excerpts from 25 English literary works, and 3) 50 custom-authored drama topics. For the HAMLET Leaderboard evaluation result in Table 1, the score for both English and Chinese is calculated on a 75-item set, which consists of the 25 literary excerpts and 50 custom topics in the corresponding language.
@@
99. James Bond and Q test new gadgets on a secret mission in Monaco.
100. Alice falls down the rabbit hole again, this time meeting characters from multiple literary worlds in Wonderland.
Appendix B B. Scoring Guideline for HAMLETJudge
HAMLETJudge which we use for critic and scoring employs the following 5-point Likert scale for pairwise comparisons. In prompt context, Model A refers to the model being evaluated, and Model B refers to the baseline model. The specific definition for each score is detailed in Table 5.
@@
Appendix C C. Case Study
@@
Table 6 presents representative real-time interaction cases generated by the HAMLET framework, showcasing the distinct roles of individual agents and their collaborative behavior during online performance. Through these cases, we illustrate how HAMLET maintains narrative coherence, ensures logical consistency, and enables flexible dramatic progression in complex and unpredictable user interactions.
@@
To begin with, Case 1 highlights the Narrator’s ability to interpret user intent under ambiguity. Specifically, it correctly associates the term “knife” with the existing prop “dagger,” allowing the user’s action to be executed successfully. This demonstrates HAMLET’s robustness in resolving lexical variations and grounding user input in the current scene context.
@@
Building on this, Cases 2,3 and 4 involve human user role-playing as AI actor, and reveal how the system handles a wide range of irregular or disruptive inputs. In Case 2, the Narrator identifies an inappropriate and non-existent prop (“rifle”), and rejects the action to preserve setting consistency. Similarly, Case 3 showcases how physically impossible actions, such as “flying,” are filtered out based on the play’s realistic constraints. Notably, Case 4 presents a more complex situation, where the user repeatedly insists on an invalid action. Here, HAMLET effectively coordinates among multiple agents: the Narrator detects the invalid target, the Transfer identifies unmet flags, and the Advancer eventually intervenes with adaptive guidance to resolve narrative deadlock. This case illustrates the system’s ability to detect stagnation, maintain progression, and deliver context-sensitive instructions.
@@
In contrast to the above, Cases 5 and 6 focus on the role of the Planner, demonstrating HAMLET’s support for multi-trajectory story planning. Both cases revolve around solving a murder mystery, yet they employ different investigative strategies. In Case 5, the AI actor (Sherlock Holmes) uncovers clues through physical evidence at the crime scene. In Case 6, the same conclusion is reached via testimonial interrogation of other characters. Despite following distinct paths, both trajectories lead to the same dramatic outcome and are validated by the Planner based on their internal logic and consistency. These cases underscore the system’s flexibility in allowing diverse narrative flows, provided the progression remains coherent and goal-aligned.
@@
Taken together, these case studies demonstrate the effectiveness of HAMLET in managing interactive drama through robust input interpretation, constraint enforcement, multi-agent collaboration, and trajectory-level planning.
@@ 4 Human Player
Butler: ”I—I was inventorying wine in the kitchen, sir.”
@@
Maid: (Wringing her apron) ”He and Mr. Bates—the butler—had shouting matches all week… something about money…”
@@
Business Partner: (Forced laugh) ”Oh, he loved boasting—said he ’had Bates by the throat’ last Tuesday.”
@@
Holmes: (Striding to the center of the room) The butler’s missing cufflink places him at the crime scene. The maid confirms violent quarrels over money, while our ’helpful’ friend here reveals the victim’s blackmail scheme. Motive, opportunity, and physical evidence—all point to one man.”
@@
Table 6: Case study of HAMLET that illustrates the roles and collaborative processes of each agent during online performance.
Appendix D D. Detailed explanation of the operation process of Perceive and Decide Module
This section provides a detailed breakdown of the operational workflow for PAD. The entire process is formally presented in Algorithm 2.
The model operates by first perceiving the complete dramatic context. This includes the current scenario $\mathcal{S}$, the actor profiles $\mathbb{C}$ and typically goals of all actors $\mathbb{G}_{c}$ defined in profile $\mathbb{C}$, the set of interactable objects $\mathbb{O}$, and dialogue history $\mathcal{H}_{d}$. Based on this comprehensive input, PAD then decides the most appropriate tool calls for each relevant actor. This decision manifests as two possible outputs: a speech pace $r$ and a potential physical action $o$, formulated as a Subject-Verb-Object triplet. This perceive-decide cycle repeats, enabling dynamic and context-aware character behaviors.
@@
41:     return $\arg\max_{r\in\{\texttt{FAST},\texttt{SLOW},\texttt{SILENCE}\}}P(r|\hat{r})$
@@
48:         $v\leftarrow\texttt{ExtractVerb}(\hat{o})$
49:         $o\leftarrow\texttt{ExtractObject}(\hat{o})$
@@
Appendix E E. Labeling Instruction for HAMLETJudge
@@ E.2 Evaluation Dimensions
  • Consistency: Are the characters’ actions and dialogues consistent with their established personalities, motivations, and backgrounds? Do they act ”out-of-character”?
@@
This dimension evaluates the overall flow, pacing, and immersive quality of the drama—essentially, its ”readability” and ”feel.” When evaluating IE, consider:
@@ E.4 Example
Detective Smith entered the dark exhibition hall, turning on his flashlight. The beam cut across an empty pedestal. ”It was stolen from here,” he muttered. On the floor, he saw a faint footprint and knelt to inspect it. ”About a size 9,” he noted. He then went to question the night guard, who claimed he didn’t hear a thing all night.
@@
Upon entering the hall, Detective Smith’s first glance wasn’t at the empty pedestal, but up at the ceiling’s security camera. ”Odd,” he murmured, ”the camera was deactivated last night, but the log shows an internal authorization.” Tucked beneath the pedestal, he found a small, folded note with a riddle: ”I sing where there is light, but die where there is silence.”
@@ E.5 Important Considerations
Appendix F F. Annotation Guidelines for PAD
@@ F.1 Task Overview
The PAD model serves as the ”brain” for each AI actor, determining the underlying strategy of a response before the final words are spoken. It is inspired by human cognition, choosing between a fast, intuitive reaction (System 1) and a slow, deliberate one (System 2), with the additional option of strategic silence.
@@
  • Environment Description
  • Current Actor List
  • Dialogure History
  • Interactable Objects
@@ F.2 Three Responding Strategies: Description and Definition
Fast (Intuitive Reaction)

### v2

@@ 5.2 Case Study
To further demonstrate the working mechanism of the 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: framework and each of its component designs, we select some real cases as shown in Table 6 in Appendix C. In Case 1, the narrator first considers the situation before responding to make a reasonable judgment about the AI actor’s input. Our framework also supports users taking on a certain character role (Cases 2, 3, and 4), where the narrator, transfer, and advancer can handle various circumstances (such as non-existent props, irresponsible actions, and stubborn choices) to ensure the drama runs smoothly. On the other hand, the planner is responsible for multi-trajectory design and review: the same goal can be met by different beat trajectories as long as the process is reasonable (Cases 5 and 6).
6 Conclusion
@@
In this paper, we introduced 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, a multi-agent framework designed to address key challenges in AI-driven theatre. 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: generates a guiding narrative blueprint from a simple topic and provides agents with a perceive and decide module, enabling autonomous thinking and physical environmental interaction. To assess performance, we established a comprehensive evaluation method and a leaderboard, where our specialized critic model, 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T0.09804 0.09412 0.23137J0.09804 0.09412 0.23137u0.09804 0.09412 0.23137d0.09804 0.09412 0.23137g0.09804 0.09412 0.23137e\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, achieved top-ranking results. Extensive experiments show that our approach creates expressive and coherent theatrical experiences, paving a new path toward autonomous and immersive interactive drama.
7 Ethics Statement
This work relies on human annotation, which is critical for both the training and evaluation of our proposed models. The annotation of 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: and 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T0.09804 0.09412 0.23137J0.09804 0.09412 0.23137u0.09804 0.09412 0.23137d0.09804 0.09412 0.23137g0.09804 0.09412 0.23137e\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: datasets was conducted by a team of five human experts, each with at least one year of professional experience in relevant fields such as creative writing or drama creation. To ensure high-quality annotations, we provided clear and task-specific instructions to all labelers. In addition, the datasets underwent rigorous safety checks to ensure no personally identifiable information or potentially problematic content is available. The labeling instructions can be found in Appendix E and F.
Appendix A Detailed information of 100 Cases Selected for Public Dataset
@@
This appendix details the public dataset of 100 cases in Table 4. The dataset comprises three parts: 1) excerpts from 25 Chinese literary works, 2) excerpts from 25 English literary works, and 3) 50 custom-authored drama topics. For the 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: Leaderboard evaluation result in Table 1, the score for both English and Chinese is calculated on a 75-item set, which consists of the 25 literary excerpts and 50 custom topics in the corresponding language.
@@
Appendix B Scoring Guideline for 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T0.09804 0.09412 0.23137J0.09804 0.09412 0.23137u0.09804 0.09412 0.23137d0.09804 0.09412 0.23137g0.09804 0.09412 0.23137e\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:
0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T0.09804 0.09412 0.23137J0.09804 0.09412 0.23137u0.09804 0.09412 0.23137d0.09804 0.09412 0.23137g0.09804 0.09412 0.23137e\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: which we use for critic and scoring employs the following 5-point Likert scale for pairwise comparisons. In prompt context, Model A refers to the model being evaluated, and Model B refers to the baseline model. The specific definition for each score is detailed in Table 5.
@@
Appendix C Case Study
@@
Table 6 presents representative real-time interaction cases generated by the 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: framework, showcasing the distinct roles of individual agents and their collaborative behavior during online performance. Through these cases, we illustrate how 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: maintains narrative coherence, ensures logical consistency, and enables flexible dramatic progression in complex and unpredictable user interactions.
@@
To begin with, Case 1 highlights the Narrator’s ability to interpret user intent under ambiguity. Specifically, it correctly associates the term “knife” with the existing prop “dagger,” allowing the user’s action to be executed successfully. This demonstrates 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: ’s robustness in resolving lexical variations and grounding user input in the current scene context.
@@
Building on this, Cases 2,3 and 4 involve human user role-playing as AI actor, and reveal how the system handles a wide range of irregular or disruptive inputs. In Case 2, the Narrator identifies an inappropriate and non-existent prop (“rifle”), and rejects the action to preserve setting consistency. Similarly, Case 3 showcases how physically impossible actions, such as “flying,” are filtered out based on the play’s realistic constraints. Notably, Case 4 presents a more complex situation, where the user repeatedly insists on an invalid action. Here, 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: effectively coordinates among multiple agents: the Narrator detects the invalid target, the Transfer identifies unmet flags, and the Advancer eventually intervenes with adaptive guidance to resolve narrative deadlock. This case illustrates the system’s ability to detect stagnation, maintain progression, and deliver context-sensitive instructions.
@@
In contrast to the above, Cases 5 and 6 focus on the role of the Planner, demonstrating 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: ’s support for multi-trajectory story planning. Both cases revolve around solving a murder mystery, yet they employ different investigative strategies. In Case 5, the AI actor (Sherlock Holmes) uncovers clues through physical evidence at the crime scene. In Case 6, the same conclusion is reached via testimonial interrogation of other characters. Despite following distinct paths, both trajectories lead to the same dramatic outcome and are validated by the Planner based on their internal logic and consistency. These cases underscore the system’s flexibility in allowing diverse narrative flows, provided the progression remains coherent and goal-aligned.
@@
Taken together, these case studies demonstrate the effectiveness of 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: in managing interactive drama through robust input interpretation, constraint enforcement, multi-agent collaboration, and trajectory-level planning.
@@ 4 Human Player
Butler: “I—I was inventorying wine in the kitchen, sir.”
@@
Maid: (Wringing her apron) “He and Mr. Bates—the butler—had shouting matches all week… something about money…”
@@
Business Partner: (Forced laugh) “Oh, he loved boasting—said he ’had Bates by the throat’ last Tuesday.”
@@
Holmes: (Striding to the center of the room) The butler’s missing cufflink places him at the crime scene. The maid confirms violent quarrels over money, while our ’helpful’ friend here reveals the victim’s blackmail scheme. Motive, opportunity, and physical evidence—all point to one man.
@@
Table 6: Case study of 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: that illustrates the roles and collaborative processes of each agent during online performance.
Appendix D Detailed explanation of the operation process of Perceive and Decide Module
This section provides a detailed breakdown of the operational workflow for 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:. The entire process is formally presented in Algorithm 2.
The model operates by first perceiving the complete dramatic context. This includes the current scenario $\mathcal{S}$, the actor profiles $\mathbb{C}$ and typically goals of all actors $\mathbb{G}_{c}$ defined in profile $\mathbb{C}$, the set of interactable objects $\mathbb{O}$, and dialogue history $\mathcal{H}_{d}$. Based on this comprehensive input, 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: then decides the most appropriate tool calls for each relevant actor. This decision manifests as two possible outputs: a speech pace $r$ and a potential physical action $o$, formulated as a Subject-Verb-Object triplet. This perceive-decide cycle repeats, enabling dynamic and context-aware character behaviors.
@@ Appendix D Detailed explanation of the operation process of Perceive and Decide Module
[…]
@@
[…]
@@
Appendix E Labeling Instruction for 0.09804 0.09412 0.23137H0.09804 0.09412 0.23137A0.09804 0.09412 0.23137M0.09804 0.09412 0.23137L0.09804 0.09412 0.23137E0.09804 0.09412 0.23137T0.09804 0.09412 0.23137J0.09804 0.09412 0.23137u0.09804 0.09412 0.23137d0.09804 0.09412 0.23137g0.09804 0.09412 0.23137e\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:
@@ E.2 Evaluation Dimensions
  • Consistency: Are the characters’ actions and dialogues consistent with their established personalities, motivations, and backgrounds? Do they act “out-of-character”?
@@
This dimension evaluates the overall flow, pacing, and immersive quality of the drama—essentially, its “readability” and “feel.” When evaluating IE, consider:
@@ E.4 Example
Detective Smith entered the dark exhibition hall, turning on his flashlight. The beam cut across an empty pedestal. “It was stolen from here,” he muttered. On the floor, he saw a faint footprint and knelt to inspect it. “About a size 9,” he noted. He then went to question the night guard, who claimed he didn’t hear a thing all night.
@@
Upon entering the hall, Detective Smith’s first glance wasn’t at the empty pedestal, but up at the ceiling’s security camera. “Odd,” he murmured, “the camera was deactivated last night, but the log shows an internal authorization.” Tucked beneath the pedestal, he found a small, folded note with a riddle: “I sing where there is light, but die where there is silence.”
@@ E.5 Important Considerations
Appendix F Annotation Guidelines for 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:
@@ F.1 Task Overview
The 0.09804 0.09412 0.23137P0.09804 0.09412 0.23137A0.09804 0.09412 0.23137D\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: model serves as the ”brain” for each AI actor, determining the underlying strategy of a response before the final words are spoken. It is inspired by human cognition, choosing between a fast, intuitive reaction (System I) and a slow, deliberate one (System II), with the additional option of strategic silence.
@@
  • Current Actor List — A real-time roster of all characters present in the scene, including their roles, relationships to the focal AI actor, and current emotional or behavioral states (e.g., “hostile,” “distracted,” “allied”). This informs social dynamics and potential audience effects on the decision-making process.
  • Dialogue History — The chronological record of all spoken (or written) exchanges up to this moment, capturing tone, unresolved tensions, promises, or prior commitments that may influence the character’s strategic choice. Subtext and implied meanings are included where relevant.
  • Interactable Objects — A catalog of physical or digital items within the environment that the character could use, reference, manipulate, or avoid. Each object may carry symbolic weight, functional utility, or emotional resonance (e.g., a flickering lantern, a locked diary, a ringing phone).
@@ F.2 Three Responding Strategies: Description and Definition
F.2.1 Fast (Intuitive Reaction)